OpenStudy (anonymous):

Kelsey painted 1/4 of her bedroom in 3/7 of an hour. At this rate, how long would it take her to paint the entire room?

OpenStudy (mathstudent55):

You can use a proportion. To do a complete job is 1. She did 1/4 of the job, so she did 1/4. 1/4 job is to 3/7 hour as 1 job is to x hours. 1/4 : 3/7 = 1 : x \(\dfrac{\frac{1}{4}}{\frac{3}{7}} = \dfrac{1}{x} \) Now cross multiply and solve for x.